X=7w​^2+4w-6
​Y=w^2-11w+13
X+Y=

Found 3 solutions by josgarithmetic, ikleyn, MathTherapy:
Answer by josgarithmetic(39623)
Answer by ikleyn(52835)

.
X + Y = 8w^2 - 7w + 7.


Answer by MathTherapy(10555)


X=7w​^2+4w-6
​Y=w^2-11w+13
X+Y=
Try adding them yourself! I'm sure you won't get an INCORRECT answer of: 8w^2 - 7w + 19 (WRONG, WRONG, WRONG)!!.